Two billiard balls of equal mass undergo an elasti
Answer #1

we will use the conservation of momnetum in x-y direction,

0 + 0 = mv sin 37 - mVsin53-------------along y axis

vsin37 = Vsin 53

v= V(0.798)/ ( 0.602)

v = 1.325V apprx-------eq(1)

m(12) + m(0) = m( v) c0s 37 + m(V) cos 53-------------along x axis

12m = m( vcos37 + Vcos53)

12= v(0.798) + V( 0.602)

12 = 1.325V (0.798) + V ( 0.602)

12= 1.66 V

V = 12/ 1.66= 7,23 m/s apprx

v = 9.578 m/s apprx